OpenStudy (anonymous):

A football is kicked toward the goal. The height of the ball is modeled by the function h(x) = -16t^2 + 64t where t equals the time in seconds and h(x) represents the height of the ball at time t seconds. What is the axis of symmetry and what does it A.x = 2; it takes 2 seconds to reach the maximum height and 2 seconds to fall back to the ground B.x = 2; it takes 2 seconds to reach the maximum height and 4 seconds to fall back to the ground C. x = 4; it takes 4 seconds to reach the maximum height and 4 seconds to fall back to the ground D. x = 4; it takes 4 seconds to reach the maxim

OpenStudy (anonymous):

Anybody?

OpenStudy (anonymous):

Well you have the function \(h(x)=-16t^2+64t=-16(t^2-4t)=-16(t)(t-4)\) From there we know that at t=4, the height is zero. Also we know that the maximum can be found by maximizing the function: \[-16t^2+64t=-16(t^2-4t)=-16(t^2-4t+4-4)=-16(t^2-4t+4)+64\] \[-16(t^2-4t+4)+64=-16(t-2)^2+64\] This means the axis of symmetry is at t=2

OpenStudy (anonymous):

is it a?

OpenStudy (anonymous):

@KeithAfasCalcLover

OpenStudy (anonymous):

Yes :)